OpenStudy (anonymous):

Write the equation so that y is a factor of x : 14=7y-6x . Please help & EXPLAIN!

OpenStudy (amistre64):

use algebra to get y all by its little lonesome

OpenStudy (amistre64):

add off the 6x and divide out the 7

OpenStudy (amistre64):

i get this when I do that: y = (6/7)x + 2

OpenStudy (amistre64):

if a read it right that is

OpenStudy (amistre64):

y is a factor of x just reads odd

OpenStudy (anonymous):

or to be a little more clear: 14 = 7y - 6x -14 -14 0 = 7y - 6x - 14 -7y -7y -7y = -6x - 14 mutiply times -1 (both left and right) 7y = 6x + 14 /7 /7 y = (6/7)x + 2 Now y is a factor of x (you can read y is a function of x)